OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x + 5 > 9 and 5x less than or equal to 30 Answer choices: A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right. A number line with an open circle on 2, a closed circle on 6, and shading in between. A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right. A number line with a closed circle on 2, an open circle on 6, and shading in between.

OpenStudy (anonymous):

Start by simplifying each statement.

OpenStudy (anonymous):

(each of the two inequalities, I mean. The one given in algebra and the other given in words.)

OpenStudy (anonymous):

I don't know how to do that,

OpenStudy (anonymous):

@hartnn @Preetha ??

hartnn (hartnn):

2x + 5 > 9 subtract 5 from both sides, what you get ?

OpenStudy (anonymous):

2x > 4

hartnn (hartnn):

yes, so divide 2 from both sides....to isolate 'x'

OpenStudy (anonymous):

x > 2 ?

hartnn (hartnn):

yes, thats correct. now 2nd part 5x less than or equal to 30 \(5x \le 30\) can you isolate x here ?

OpenStudy (anonymous):

\[x \le 6 ?\]

hartnn (hartnn):

yes, so now you have to simplify , \(x>2 \quad and \quad x\le 6\)

hartnn (hartnn):

can you ?

hartnn (hartnn):

or you can look at the options and figure out

OpenStudy (anonymous):

I don't know how to do that.

hartnn (hartnn):

x>2 means there will be shading on right of 2 since x cannot be =2, the circle will be open,  got this part ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

x<= 6 means shading to left of 6 and since x can take value of 6, its closed circle,  so now what you think ? which option ?